OpenStudy (anonymous):

Medal to whoever helps. (Algebra 1)

geerky42 (geerky42):

That's one long username @ohlookidontunderstandanotherthing

OpenStudy (anonymous):

I can try to help!! Hello person from The attack on titans fandom!

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Hello Whovian

OpenStudy (anonymous):

I know. I don't know what I was thinking when I made it my name @geerky42

OpenStudy (anonymous):

It's unique. :) @ohlookidontunderstandanotherthing

OpenStudy (anonymous):

Thanks :D @Prabjotp12

OpenStudy (anonymous):

ok so here is the question The table below represents the distance a person runs as a function of time: The average rate of change of the function between x = 10 to x = 16 is __________ m/s and represents the person's speed.

OpenStudy (anonymous):

here is the table Time Distance (x) f(x) 10 60 13 78 16 96 19 114

OpenStudy (anonymous):

the equation thing to find rate of change is.\[\frac{ f(b) - f(a )}{ b-a }\]

OpenStudy (anonymous):

your welcome! @ohlookidontunderstandanotherthing

OpenStudy (anonymous):

I assume the denominator is for X's? So if my assumption is correct then your equation shall be \[\frac{ 96-60 }{ 16-19 }\]

OpenStudy (anonymous):

Can you solve the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

\[\frac{ 36 }{ 6 }\]

OpenStudy (anonymous):

should I simplify it?

OpenStudy (anonymous):

Yes! Correct! You can simplify it! I had my equation wrong with the 9 inlace of the 10 but you did it correctly. Nice job!

OpenStudy (anonymous):

so it should end up being 6?

OpenStudy (anonymous):

Yes! Nice job!

OpenStudy (anonymous):

can you help me with another one?

OpenStudy (anonymous):

I can try!!

OpenStudy (anonymous):

Awesome! The following function defines a recursive sequence: f(0) = -5 f(1) = 20 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? a. -5, -20, -65, -200, … b. -5, 20, -92, 372, … c. -5, -24, -92, -372, … d. -5, 20, -65, 200, …

OpenStudy (anonymous):

This one I have no clue. Sorry! I never really grasped the recursive sequences. I'd much rather someone with a better understanding give you the correct answer then me giving you the wrong answer. Good luck!

OpenStudy (anonymous):

OK thanks anyways you've been a big help!